OpenStudy (anonymous):

Write an equation in point-slope form of the line that passes through the points (-3,-4) and (3,4).

OpenStudy (tylerd):

how would you find the slope?

OpenStudy (tylerd):

\[\frac{ 4-(-4) }{ 3-(-3) }\]

OpenStudy (tylerd):

\[M=\frac{ 8 }{ 6 }\]

OpenStudy (anonymous):

soo now its y=8/6x-or +b rite Tyler

OpenStudy (tylerd):

Y=8/6x+b yep

OpenStudy (tylerd):

in this case i think B is 0

OpenStudy (tylerd):

wait

OpenStudy (tylerd):

yep thats it y=8/6x

OpenStudy (tylerd):

it intersects the origin, we can tell by the (-3,-4) (3,4) same numbers except one point is negative numbers.

OpenStudy (tylerd):

when the y intercept is 0, B=0 in this case.

OpenStudy (anonymous):

so tha answer is y=8/6x+b rite

OpenStudy (tylerd):

b = 0 so there is no b

OpenStudy (tylerd):

y=8/6x+0 minus well just write y=8/6x

OpenStudy (tylerd):

study the material though because your teacher may ask how you get the answers.

OpenStudy (anonymous):

oh ok cool now I have another ques

OpenStudy (tylerd):

ive gotta go, sorry.

OpenStudy (anonymous):

oh okay do u know any1 else who can help